Lauren has $1000 made up of only ten and twenty dollar bills. She has a total of 75 bills. How many of each kind does she have?
Answer by KnightOwlTutor(293)

Let # of $10=x
Let# of $20=75-x
Here is the equation
$10x+$20(75-x)=$1000
Use the distributive property

$10x+$1500-$20x=$1000
Subtract 1500 from both sides
add the two x terms
-$10x=-$500
Divide both sides by -10
x=50
Let's check our answer
when the number of $10 bills is 50 you have $10(50)=$500
75-25=50 therefore the $ amount of &20=$500
$500+$500=$1000 which is the amount of $ that Lauren has.